6. (7 pts.) The mean age of all used cars for sale in a newspaper one Saturday last month was7 years, with a standard deviation of 4 years. The distribution of ages is right-skewed. For a study to determine the reliability of classified ads, a reporter randomly selects 36 of these used (a) Mean of Distribution of T cars. SE of- (b) Find the probability that the mean age of the 49 cars selected is more than 8 years. P(E> 8)
Answer #1

Solution:

Given that,

mean = \mu = 7 years

standard deviation = \sigma = 4 years

n = 36

So,

a ) Mean distribution is  \mu\bar x

  \mu\bar x = 7

Standard error of mean distribution is  \sigma\bar x

\sigma\bar x =  (\sigma /\sqrtn)

\sigma\bar x  = ( 4 / \sqrt36 )

\sigma\bar x= 0.6667

b ) n = 49

So,

  \mu\bar x = 7

\sigma\bar x =  (\sigma /\sqrtn) = ( 4 / \sqrt49 ) = 0.5714

p (  \bar x > 8 )

= 1 - p (  \bar x < 8 )

= 1 - p ( \bar x - \mu\bar x/\sigma\bar x) < ( 8 - 7 /0.5714 )

= 1 - p ( z < 1 /0.5714 )

= 1 - p ( z < 1.75 )

Using z table

= 1 - 0.9599

= 0.0401

Probability = 0.0401